# Many-Body Localization in the Age of Classical Computing

### v2

@@ IX.1.4 Open system approach
The difficulties in finding direct numerical evidence pin-pointing the mechanism of the MBL transition motivate the search for alternative approaches. Ref. [427, 571] proposes to examine an open system described by the GKSL equation (66) with Lindblad operators
$L_{k}=\{S^{x}_{1},S^{y}_{1},S^{z}_{1}\},$ (82)
@@
where $\lambda_{k},\rho_{k}$ are eigenvalues and eigenvectors of the Liouvillian superoperator $\hat{\cal L}$ defined in (66), i.e., ^$\hat{\cal L}(\rho_{k})=\lambda_{k}\rho_{k}$. For the chosen Lindblad operators, the steady state $\rho_{\infty}=\lim_{t\to\infty}\rho(t)$ is unique and corresponds to an infinite temperature state. At the same time, the slowest decaying mode is associated with the eigenvalue $\lambda_{1}$ with the second largest real part; its relaxation rate is given by $\Gamma=-\Re(\lambda_{1})$.
@@
Ref. [427, 571] propose that insights into the stability of MBL in a disordered spin chain coupled to a large thermal inclusion can be obtained by the analysis of the GKSL equation with Lindblad operators $L_{k}$ given by Eq. (82). According to this proposition, the MBL is stable to avalanches spreading throughout the system if the relaxation rate of the slowest decaying mode, $\Gamma$, decreases with system size $L$ faster than $4^{-L}$. This condition, in the perturbative limit of weak coupling to the Markovian bath described by (66), is fulfilled by the disordered XXZ spin chain at $W^{\mathrm{avch}}(L=14)\approx 20$ [571], exceeding the value of $W^{*}(L=14)$ extracted from the standard indicators of ergodicity breaking available in exact diagonalization studies by a factor of more than $5$. The disorder strength $W^{\mathrm{avch}}(L)$ is compared by [427] to other disorder strengths which are sensitive to the emergence of resonances between different spin configurations in eigenstates of (7). The latter disorder strengths exceed the value of $W^{*}(L)$ by a factor of $2$ but are still well below the value of $W^{\mathrm{avch}}(L)$. We revisit their relation in Sec. IX.2.3.
@@
It is worth emphasizing that the connection between a GKSL equation with the Lindblad operators (82) and the spreading of avalanches seeded by an ergodic inclusion needs to be further clarified. In particular, it is even not apparent whether a large ergodic inclusion formed by a weakly disordered region in the XXZ spin chain fulfills the stringent assumptions behind the GKSL equation discussed here. One possible reconciliation of the value of $W^{\mathrm{avch}}(L)$ with the exact diagonalization results, which place the ETH-MBL crossover at significantly smaller disorder strengths, is that the effects of rare large thermal regions are absent at system sizes accessible to exact diagonalization [571]. This perspective is, however, discouraged by the results of [634]. Despite the planted thermal inclusions are likely to be found only in systems as large as $L\approx 10^{6}$[634] do not observe any significant drift of the disorder strengths at which qualitative changes between the ergodic and MBL time evolutions are observed. Therefore, further investigations of microscopic mechanisms behind the avalanche propagation [257] and the emergence of ergodicity in disordered spin chains are needed. Such studies should also consider the more complicated types of dynamics allowing for intermediate dynamical regimes [144, 145].
@@ IX.2.1 Thermalization of dilute impurities
The idea of the avalanche scenario is to examine the stability of LIOMs to rare thermal regions. Let us assume that the conditions for the propagation of avalanches are fulfilled. In that case, the MBL regions are destabilized due to the coupling to rare thermal regions, and the whole system becomes ergodic. Ref. [573] proposes to reverse this perspective and look for a mechanism for MBL transition by examining the fate of a finite density of strongly disordered sites, referred to as "impurities," in an otherwise ergodic system.
@@
Focusing on a single impurity case, when the set of the sites with impurities contains a single element, =${\cal I}=\{(L+1)/2\}$, Ref. [573] demonstrates that the impurity is associated with a LIOM (which is adiabatically linked to the $S^{z}_{(L+1)/2}$ operator in the $V\to\infty$ limit) provided that the system size is smaller than a threshold value which scales linearly with the impurity potential strength $V$ up to logarithmic corrections. Hence, the system is globally ergodic for $V$ below $V^{*}\propto L$. Consequently, the average gap ratio $\overline{r}$ starts deviating from the GOE value at $V^{*}\propto L$, marking the onset of delocalization of the LIOM associated with the impurity. The subsequent increase of $V$ beyond $V^{*}$ leads to a further decrease of the value of $\overline{r}$, as the parts of the chain on the two sides of the impurity are connected at large $V$ only by the virtual terms $\propto 1/V$.
@@
Ref. [573] also considers the case of a finite density of impurities. In particular, by placing the impurities on every fourth lattice site, the impurity potential strength $V^{*}$, at which the deviation of $\overline{r}$ from the GOE value is visible, is again scaling linearly with system size, $V^{*}\propto L$ (within the examined interval of system sizes $L<20$). However, an increasing number of lattice sites, associated with increase of the number of impurities in the system, leads to a growth of the number of sectors into which the chain gets broken in the large $V$ limit. The increased number of blocks into which the Hilbert space is effectively fragmented causes an additional decrease of $\overline{r}$ at $V>V^{*}$ towards the Poisson value. This reduction of the value of $\overline{r}$ at $V>V^{*}$ leads to an appearance of crossing points of the $\overline{r}(V)$ curves. However, these crossing points do not convey information about the ergodicity breakdown in the system, as the LIOMs are already destabilized at the weaker impurity potential $V=V^{*}\propto L$.
@@
These observations, substantiated by insights into the behavior of the spectral functions and operator spreading in the system, cause the authors of Ref. [573] to conclude that the system remains ergodic for finite impurity densities at system sizes beyond $L\propto V^{*}$. The appearance of the crossing points for the $\overline{r}$ is a result of the fragmentation of the Hilbert space into an increasing number of sectors rather than a signal of the breakdown of the ergodicity. The breakdown of the ergodicity occurs already at $V^{*}\propto L$. The latter behavior is analogous to the linear drift $W^{T}(L)\propto L$ found in the results for the disordered spin chains as discussed in Sec. VI.1.1. If these linear system size drifts persist to $L\to\infty$, the spin chains are ultimately, in the thermodynamic limit, ergodic at any $V$. Regardless of whether that is indeed the case, the impurity approach enhances our understanding of finite-size drifts in disordered spin chains, providing a helpful perspective on MBL transition. Yet, in specific Floquet systems, the linear drift $W^{T}(L)\propto L$ is broken at sufficiently large system sizes [596]. Therefore, computational studies of linear drifts at large system sizes in autonomous disordered spin chains may shed further light on the status of MBL.
IX.2.2 A model of many-body resonances
@@
Broadly speaking, many-body resonances occur whenever an eigenstate of a many-body system can be expressed as a superposition of a few weakly entangled states, e.g., the computational basis states, that substantially differ in extensively many local regions. Dynamically, the many-body resonances manifest as large oscillations in correlation functions of local observables.
@@
where the correlation length $\xi$ sets the typical range of resonances, while the resonance length $\lambda$ determines the density of resonances. The approach of [146] employs perturbative arguments to obtain insights into the behavior of $\xi$ and $\lambda$ in disordered spin chains. These arguments suggest that $\xi$ diverges at the MBL transition and that the finite-size behavior in the vicinity of the transition is determined by the resonance length $\lambda$. Notably, the value of the latter parameter corresponding to the physics of the disordered XXZ spin chain is estimated as $15\lessapprox\lambda\lessapprox 50$. For that parameter, the resonance model of [146] reproduces many features of the results observed numerically for disordered spin chains, including: the linear drift of the disorder strength $W^{T}(L)$ at the ETH-MBL crossover, slow dynamics, and the apparent $1/\omega$ behavior of the spectral function in the crossover regime, or exponential increase of Thouless time $t_{Th}$ at small disorder strengths. Notably, all of those features are obtained within the resonance model, which, by construction, hosts an MBL phase. While the resonance model of [146] relies on the first-order perturbation theory, which becomes invalid in the vicinity of the ergodic phase, it provides the basis for interpretation of the phenomenology of prethermal MBL proposed in [380], as discussed in Sec. VII.3.2.
The model of many-body resonances considered in [219] employs the perspective of parametric level dynamics induced by smooth variations of the disorder, see Sec. VI.3.2, to understand the properties resonances between the LIOM configurations in the MBL regime. This framework allowed to determine the distributions of off-diagonal matrix elements of local operators and their correlations with the level spectrum. A subsequent work [220] provided an extension of this framework, formulated without a direct reference to the picture of LIOMs. Ref. [682] used numerically constructed LIOMs to investigate the many-body resonances, finding hybridizations at all considered length scales in the vicinity of the ETH-MBL crossover.
@@ IX.2.3 Probing many-body resonances with exact diagonalization
In addition to the open system approach and the corresponding disorder strength $W^{\mathrm{avch}}(L)$, Ref. [427] proposes several probes of many-body resonances in eigenstates of disordered spin chains.
Quantum mutual information (QMI) $I^{\mathrm{e-e}}_{2}$ (39) between the ends of the chain (assuming open boundary conditions), i.e., for subsystems $A_{1}=\{1\}$ and $A_{2}=\{L\}$ corresponding to the first ($j=1$) and the last ($j=L$) spin, allows to examine emergence of many-body resonances. Deep in the ergodic phase, the QMI $I^{\mathrm{e-e}}_{2}$ is small since the entanglement entropy follows the volume-law scaling and the contribution $S_{A_{1}}+S_{A_{2}}$ approximately cancels out with $S_{A_{1}\cup A_{2}}$. Similarly, the QMI vanishes deep in the MBL regime when the end spins are not correlated. In contrast, $I^{\mathrm{e-e}}_{2}$ is enhanced in the intermediate crossover regime between the ergodic and MBL regimes when the eigenstates are superpositions of a limited number of basis states with substantially different spin configurations. For that reason, $I^{\mathrm{e-e}}_{2}$ is well suited for investigation of system-wide resonances appearing whenever an eigenstate is a Schrödinger cat-like state composed of a few basis states. To probe the onset of such resonances, [427] consider the maximum $I^{\mathrm{e-e}}_{\mathrm{max}}$ of $I^{\mathrm{e-e}}_{2}$ over all eigenstates of the disordered XXZ spin chain, and take the median of $I^{\mathrm{e-e}}_{\mathrm{max}}$ over disorder realizations.
The resulting quantity, plotted as a function of $W$ for various $L$ has a crossing point $W^{\mathrm{swr}}(L)$, similarly to $\overline{r}$ shown in Fig. 4. Notably, the crossing points $W^{\mathrm{swr}}(L)$ occur at significantly larger disorder strengths than $W^{*}(L)$. For instance, for disordered XXZ spin chain, $W^{*}(L=16)\approx 3.5$, while $W^{\mathrm{swr}}(L=16)\approx 8.5$. The drift of $W^{\mathrm{swr}}(L)$ with an increase of system size $L$, relative to its value at small system sizes, is comparable to the drift of $W^{*}(L)$. In particular, between $L=11$ and $L=16$ the value of $W^{\mathrm{swr}}(L)$ increases by approximately $1/3$ of its value. Notably, however, the value of the QMI $I^{\mathrm{e-e}}_{\mathrm{max}}$ at the crossing point $W=W^{\mathrm{swr}}(L)$ is of the order of $10^{-7}$. Hence, the resonances found by this quantity are very weak. On the other hand, these resonances extend between the ends of the whole chain, indicating some form of breakdown of locality in the system.
@@
Juxtaposing the values of $W^{\mathrm{swr}}(L)$, $W^{\mathrm{mg}}(L)$ and $W^{\mathrm{avch}}(L)$ with the value of $W^{*}(L)$, Ref. [427, 571] conclude that MBL phase, if it exists at all in the XXZ spin chain, can be present only at disorder strengths larger than $W\gtrapprox 20$. These results lead to two critical questions which remain open. The first question is whether $W^{\mathrm{avch}}(L)$, obtained in the open system approach, is relevant for the physics of isolated disordered spin chains. The second question is about understanding the relation of $W^{\mathrm{swr}}(L)$ and $W^{\mathrm{mg}}(L)$, which are obtained as the extreme values of ergodicity breaking indicators, with the standard quantities analyzed in the context of ETH-MBL crossover such as $W^{*}(L)$. We can improve our understanding of this connection through computational studies of disordered spin chains and reference systems, such as the Anderson model on random regular graphs, or random matrix models [348], even at system sizes available for present-day classical computers. An investigation of extreme value statistics of an average of local observable in disordered XXZ spin chains reports a sharp change of the properties of extreme value statistics at $W=4-7$ [135]. These avenues constitute intriguing directions for further studies of MBL and may shed new light on the existence and critical disorder strength of the MBL transition.

### v3

@@ IX.1.4 Open system approach
The difficulties in finding direct numerical evidence pin-pointing the mechanism of the MBL transition motivate the search for alternative approaches. Ref. [463, 618] proposed to examine an open system described by the GKSL equation (66), with the Lindblad operators
$\hat{L}_{k}=\{\hat{S}^{x}_{1},\hat{S}^{y}_{1},\hat{S}^{z}_{1}\},$ (82)
@@
where $\lambda_{k},\hat{\rho}_{k}$ are the eigenvalues and right eigenvectors of the Liouvillian superoperator $\hat{\cal L}$ defined in (66), i.e., ^$\hat{\cal L}(\hat{\rho}_{k})=\lambda_{k}\hat{\rho}_{k}$, and the coefficients $p_{k}$ are the overlaps of the initial state with the left eigenvectors of $\hat{\cal L}$. For the chosen Lindblad operators, the steady state $\hat{\rho}_{\infty}=\lim_{t\to\infty}\hat{\rho}(t)$ is unique and corresponds to an infinite temperature state, $\hat{\rho}_{\infty}\propto\mathbb{I}$, corresponding to the eigenvalue $\lambda_{0}=0$. At the same time, the slowest decaying mode is associated with the eigenvalue $\lambda_{1}$ with the second largest real part; its relaxation rate is given by $\Gamma=-\Re(\lambda_{1})$.
@@
Ref. [463, 618] proposed that insights into the stability of MBL in a disordered spin chain coupled to a large thermal inclusion can be obtained by the analysis of the GKSL equation with Lindblad operators $\hat{L}_{k}$ given by Eq. (82). According to this proposition, the MBL is stable to avalanches spreading throughout the system if the relaxation rate of the slowest decaying mode, $\Gamma$, decreases with system size $L$ faster than $4^{-L}$. This condition, in the perturbative limit of weak coupling to the Markovian bath described by (66), is fulfilled by the disordered XXZ spin chain at $W^{\mathrm{avch}}(L=14)\approx 20$ [618], exceeding the value of $W^{*}(L=14)$ extracted from the standard indicators of ergodicity breaking available in exact diagonalization studies by a factor of more than $5$. The disorder strength $W^{\mathrm{avch}}(L)$ is compared by [463] to other disorder strengths which are sensitive to the emergence of resonances between different spin configurations in eigenstates of (7). The latter disorder strengths exceed the value of $W^{*}(L)$ by a factor of $2$ but are still well below the value of $W^{\mathrm{avch}}(L)$. We revisit their relation in Sec. IX.2.3.
@@
It is worth emphasizing that the connection between a GKSL equation with the Lindblad operators (82), and the spreading of avalanches seeded by an ergodic inclusion, needs to be further clarified. In particular, it is even not apparent whether a large ergodic inclusion formed by a weakly disordered region in the XXZ spin chain fulfills the stringent assumptions behind the GKSL equation discussed here. One possible reconciliation of the value of $W^{\mathrm{avch}}(L)$ with the exact diagonalization results, which place the ETH-MBL crossover at significantly smaller disorder strengths, is that the effects of rare large thermal regions are absent at system sizes accessible to exact diagonalization [618]. This perspective is, however, discouraged by the results of [685]. Despite the planted thermal inclusions are likely to be found only in systems as large as $L\approx 10^{6}$[685] do not observe any significant drift of the disorder strengths at which qualitative changes between the ergodic and MBL time evolutions are observed. Therefore, further investigations of microscopic mechanisms behind the avalanche propagation [273] and the emergence of ergodicity in disordered spin chains are needed. Such studies should also consider the more complicated types of dynamics allowing for intermediate dynamical regimes [149, 150].
@@ IX.2.1 Thermalization of dilute impurities
The idea of the avalanche scenario is to examine the stability of LIOMs to rare thermal regions. Let us assume that the conditions for the propagation of avalanches are fulfilled. In that case, the MBL regions are destabilized due to the coupling to rare thermal regions, and the whole system becomes ergodic. Ref. [620] proposed to reverse this perspective and look for a mechanism of MBL transition by examining the fate of a small density of strongly disordered sites, referred to as "impurities," in an otherwise ergodic system.
@@
Focusing on a single impurity case, when the set of the sites with impurities contains a single element, =${\cal I}=\{(L+1)/2\}$, Ref. [620] demonstrated that the impurity is associated with a LIOM (which is adiabatically linked to the $\hat{S}^{z}_{(L+1)/2}$ operator in the $V\to\infty$ limit) provided that the system size is smaller than a threshold value, which scales linearly with the impurity potential strength $V$ up to logarithmic corrections. Hence, the system is globally ergodic for $V$ below $V^{*}\propto L$ and the LIOM associated with the impurity becomes delocalized at $V<V^{*}$. Consequently, the average gap ratio $\overline{r}$ starts deviating from the GOE value at $V^{*}\propto L$. The subsequent increase of $V$ beyond $V^{*}$ leads to a further decrease of the value of $\overline{r}$, as the parts of the chain on the two sides of the impurity are connected at large $V$ only by the virtual terms $\propto 1/V$.
@@
Ref. [620] also considered the case of a finite density of impurities. In particular, by placing the impurities on every fourth lattice site, the impurity potential strength $V^{*}$, at which the deviation of $\overline{r}$ from the GOE value is visible, is again scaling linearly with system size, $V^{*}\propto L$ (within the examined interval of system sizes $L<20$). However, an increasing number of lattice sites, associated with an increase of the number of impurities in the system, leads to a growth of the number of sectors into which the chain gets broken in the large $V$ limit. The increased number of blocks into which the Hilbert space is effectively fragmented causes an additional decrease of $\overline{r}$ at $V>V^{*}$ towards the Poisson value. This reduction of the value of $\overline{r}$ at $V>V^{*}$ leads to an appearance of crossing points of the $\overline{r}(V)$ curves. However, these crossing points do not convey information about the ergodicity breakdown in the system, as the LIOMs are already destabilized at the weaker impurity potential $V=V^{*}\propto L$.
@@
These observations, substantiated by insights into the behavior of the spectral functions and operator spreading in the system, caused the authors of Ref. [620] to conclude that the system remains ergodic for finite impurity densities at system sizes beyond $L\propto V^{*}$. The appearance of the crossing points for the $\overline{r}$ is a result of the fragmentation of the Hilbert space into an increasing number of sectors rather than a signal of the breakdown of the ergodicity. The breakdown of the ergodicity occurs already at $V^{*}\propto L$. The latter behavior is analogous to the linear drift $W^{T}(L)\propto L$ found in the results for the disordered spin chains as discussed in Sec. VI.1.1. If these linear system size drifts persist to $L\to\infty$, the spin chains are ultimately, in the thermodynamic limit, ergodic at any $V$. Regardless of whether that is indeed the case, the impurity approach enhances our understanding of finite-size drifts in disordered spin chains, providing a helpful perspective on MBL transition. Yet, in specific Floquet systems, the linear drift $W^{T}(L)\propto L$ is broken at sufficiently large system sizes and, instead, a sub-linear dependence of $W^{T}$ on $L$ is observed [645]. Therefore, computational studies of linear drifts at large system sizes in autonomous disordered spin chains may shed further light on the status of MBL.
IX.2.2 Models of many-body resonances
@@
Broadly speaking, many-body resonances occur whenever an eigenstate of a many-body system can be expressed as a superposition of a few not entangled or weakly entangled states, e.g., the computational basis states, that substantially differ in extensively many local regions. Dynamically, the many-body resonances manifest as large oscillations in correlation functions of local observables.
@@
where the correlation length $\xi$ sets the typical range of resonances, while the resonance length $\lambda$ determines the density of resonances. The approach of [151] employs perturbative arguments to obtain insights into the behavior of $\xi$ and $\lambda$ in disordered spin chains. These arguments suggest that $\xi$ diverges at the MBL transition and that the finite-size behavior in the vicinity of the transition is determined by the resonance length $\lambda$. Notably, the value of the latter parameter corresponding to the physics of the disordered XXZ spin chain is estimated as $15\lessapprox\lambda\lessapprox 50$. For that parameter, the resonance model of [151] reproduces many features of the results observed numerically for the disordered spin chains, including: the linear drift of the disorder strength $W^{T}(L)$ at the ETH-MBL crossover, slow dynamics, and the apparent $1/\omega$ behavior of the spectral function in the crossover regime, or the exponential increase of Thouless time $t_{Th}$ at small disorder strengths. Notably, all of those features are obtained within the resonance model, which, by construction, hosts an MBL phase. While the resonance model of [151] relies on the first-order perturbation theory, which becomes invalid in the vicinity of the ergodic phase, it provides the basis for interpretation of the phenomenology of prethermal MBL proposed in [416], as discussed in Sec. VII.3.2.
The model of many-body resonances considered in [234] employs the perspective of parametric level dynamics induced by smooth variations of the disorder, see Sec. VI.3.2, to understand the properties of resonances between the LIOM configurations in the MBL regime. This framework allowed to determine the distributions of off-diagonal matrix elements of local operators and their correlations with the level spectrum. A subsequent work [235] provided an extension of this framework, formulated without a direct reference to the picture of LIOMs. Ref. [737] used numerically constructed LIOMs to investigate the many-body resonances, finding hybridizations at all considered length scales in the vicinity of the ETH-MBL crossover.
@@ IX.2.3 Probing many-body resonances with exact diagonalization
In addition to the open system approach and the corresponding disorder strength $W^{\mathrm{avch}}(L)$, Ref. [463] proposed several probes of many-body resonances in eigenstates of disordered spin chains.
Quantum mutual information (QMI) $I^{\mathrm{e-e}}_{2}$ (39) between the ends of the chain (assuming open boundary conditions), i.e., for subsystems $A_{1}=\{1\}$ and $A_{2}=\{L\}$ corresponding to the first ($j=1$) and the last ($j=L$) spin, allows to examine emergence of many-body resonances. Deep in the ergodic phase, the QMI $I^{\mathrm{e-e}}_{2}$ is small since the entanglement entropy follows the volume-law scaling and the contribution $S_{A_{1}}+S_{A_{2}}$ approximately cancels out with $S_{A_{1}\cup A_{2}}$. Similarly, the QMI vanishes deep in the MBL regime when the end spins are not correlated. In contrast, $I^{\mathrm{e-e}}_{2}$ is enhanced in the intermediate crossover regime between the ergodic and MBL regimes when the eigenstates are superpositions of a limited number of basis states with substantially different spin configurations. For that reason, $I^{\mathrm{e-e}}_{2}$ is well suited for investigation of system-wide resonances appearing whenever an eigenstate is a Schrödinger cat-like state composed of a few basis states. To probe the onset of such resonances, [463] considered the maximum $I^{\mathrm{e-e}}_{\mathrm{max}}$ of $I^{\mathrm{e-e}}_{2}$ over all eigenstates of the disordered XXZ spin chain, and take the median of $I^{\mathrm{e-e}}_{\mathrm{max}}$ over disorder realizations.
The resulting quantity, plotted as a function of $W$ for various $L$ has a crossing point $W^{\mathrm{swr}}(L)$, similarly to $\overline{r}$ shown in Fig. 4. Notably, the crossing points $W^{\mathrm{swr}}(L)$ occur at significantly larger disorder strengths than $W^{*}(L)$. For instance, for the disordered XXZ spin chain, $W^{*}(L=16)\approx 3.5$, while $W^{\mathrm{swr}}(L=16)\approx 8.5$. The drift of $W^{\mathrm{swr}}(L)$ with an increase of system size $L$, relative to its value at small system sizes, is comparable to the drift of $W^{*}(L)$. In particular, between $L=11$ and $L=16$ the value of $W^{\mathrm{swr}}(L)$ increases by approximately $1/3$ of its value. Notably, however, the value of the QMI $I^{\mathrm{e-e}}_{\mathrm{max}}$ at the crossing point $W=W^{\mathrm{swr}}(L)$ is of the order of $10^{-7}$. Hence, the resonances found by this quantity are very weak. On the other hand, these resonances extend between the ends of the whole chain, indicating some form of breakdown of locality in the system. The maximum of $I^{\mathrm{e-e}}_{2}$ as a function of disorder strength also signals a characteristic disorder strength quantifying the MBL crossover. The maximum of $I^{\mathrm{e-e}}_{2}$ occurs at disorder strength comparable to the crossing point of gap ratio $W^{*}(L=16)\approx 3.5$ [463].
@@
Juxtaposing the values of $W^{\mathrm{swr}}(L)$, $W^{\mathrm{mg}}(L)$ and $W^{\mathrm{avch}}(L)$ with the value of $W^{*}(L)$[463, 618] concluded that the MBL phase, if it exists at all in the disordered XXZ spin chain, can be present only at disorder strengths larger than $W\gtrapprox 20$. The latter inequality stems from $W^{\mathrm{avch}}(L=14)\approx 20$, which is considerably larger than the other characteristic disorder strengths, $W^{\mathrm{swr}}(L)$, $W^{\mathrm{mg}}(L)$, and $W^{*}(L)$. This conclusion leads to two critical questions which remain open. The first question is whether $W^{\mathrm{avch}}(L)$, obtained in the open system approach, is relevant for the physics of isolated disordered spin chains. The second question is about understanding the relation of $W^{\mathrm{swr}}(L)$ and $W^{\mathrm{mg}}(L)$, which are obtained as the extreme values of ergodicity breaking indicators, with the standard quantities analyzed in the context of ETH-MBL crossover such as $W^{*}(L)$. We can improve our understanding of this connection through computational studies of disordered spin chains and reference systems, such as the Anderson model on random regular graphs, or random matrix models [377], even at system sizes available for present-day classical computers. An investigation of extreme value statistics of an average of local observable in the disordered XXZ spin chains reported a sharp change of the properties of extreme value statistics at $W=4-7$ [140]. These avenues constitute intriguing directions for further studies of MBL and may shed new light on the existence and critical disorder strength of the MBL transition.